# Reliability-Latency-Rate Tradeoff in Low-Latency Communications with Finite-Blocklength Coding

### v3

@@ Appendix F Proof of Lemma 1
$\lim_{\gamma\to+\infty}\frac{1}{\sqrt{N}}\sqrt{\frac{-\log_{2}\epsilon(\gamma)}{\log_{2}^{2}\gamma}}$ $=\lim_{\gamma\to+\infty}\frac{\sqrt{-\log_{2}\epsilon(\gamma)}}{\sqrt{N}\log_{2}\gamma}$ (151a)
$=\lim_{\gamma\to+\infty}\frac{\frac{\partial\sqrt{-\log_{2}\epsilon(\gamma)}}{\partial\log_{2}\gamma}}{\sqrt{N}+\frac{\partial\sqrt{N}}{\partial\log_{2}\gamma}\log_{2}\gamma}$ (151b)
@@
Eq. (151b) holds according to L’Hospital’s rule. As defined in Lemma 1, we let $c_{4}$ denote $\lim_{\gamma\to+\infty}\frac{1}{1+\frac{\partial\sqrt{N}}{\partial\log_{2}\gamma}\frac{\log_{2}\gamma}{\sqrt{N}}}$. Since $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$, $\frac{\partial\sqrt{N}}{\partial\log_{2}\gamma}\frac{\log_{2}\gamma}{\sqrt{N}}\geq 0$. Thus, $c_{3}\in[0,1]$.
Then, we let $Q^{-1}(\epsilon(\gamma))=x$, which can also be written as $\epsilon(\gamma)=Q(x)$. According to Eq. (102), we have
@@
By replacing $v$ by $\epsilon(\gamma)$ in Eq. (104), we have
${\color[rgb]{0,0,0}-\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}\leq x\leq\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}.}$ (153)
Since $x>0$, we have $0<x\leq\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}$. Similarly, according to Eq. (102), we obtain
Based on $W(\cdot)$, we obtain
${\color[rgb]{0,0,0}x+\frac{1}{x}\geq\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}\text{ or }x+\frac{1}{x}\leq-\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}}$ (155)
Since $x>0$, we will utilize $x+\frac{1}{x}\geq\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}$ in the following.
@@
Similarly, based on Eq. (155), we obtain
$\lim_{\gamma\to+\infty}\frac{Q^{-1}(\epsilon(\gamma))}{\sqrt{N}\log_{2}\gamma}\geq\lim_{\gamma\to+\infty}\frac{\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}}{\sqrt{N}\log_{2}\gamma}-\lim_{\gamma\to+\infty}\frac{1}{Q^{-1}(\epsilon(\gamma))\cdot\sqrt{N}\log_{2}\gamma}=\frac{c_{4}}{\log_{2}e}\varpi.$ (157)
Eq. (157) holds because with $\gamma\to+\infty$, $Q^{-1}(\epsilon(\gamma))\to+\infty$. Thus, according to the squeeze theorem, we have completed the proof.
@@ Appendix G Proof of Corollary 3
Based on Lemma 1 in [52], we find that $\alpha_{A}(\theta)$ is an increasing function of $\theta$. Thus, the inverse function of $\alpha_{A}(\cdot)$ exists, which we denoted by $\alpha^{-1}_{A}(\cdot)$. According to Eq. (20), the $\theta$ for the AWGN channel can be denoted by
@@
Based on Eq. (19), we have $\chi\leq e^{-\theta L}$. By substituting Eq. (158) into $\chi\leq e^{-\theta L}\leq\chi_{\rm th}$, we have
@@
Since $\alpha_{A}(\theta)$ is an increasing function of $\theta$, we have $R_{N}^{\epsilon}\geq\alpha_{A}\left(-\frac{1}{L}\ln\chi\right)$. By substituting Eq. (7) into this inequality, we have
${\color[rgb]{0,0,0}NC-\sqrt{NV}Q^{-1}(\epsilon)+\frac{\log_{2}N}{2}+G(N,\gamma,\epsilon)\geq\alpha_{A}\left(-\frac{1}{L}\ln\chi\right).}$ (160)
Based on Eq. (160), we obtain Eq. (70).
@@ Appendix H Proof of Lemma 2
We put our focus on the expectation term in the definition of EC. We define $g(x)$ as $-r(x)$. The derivative of $g(x)$ is given by
$(1+\gamma x)\sqrt{(1+\gamma x)^{2}-1}$ is an increasing function of $x$, which is equals to 0 with $x=0$. Thus, based on Eq. (161), we find that there exists a $x^{*}$, when $x=x^{*}$, $g^{\prime}(x)=0$, $x<x^{*}$, $g^{\prime}(x)>0$, and $x>x^{*}$, $g^{\prime}(x)<0$. $x^{*}$ satisfies
@@
$x^{*}$ can be solved from Eq. (162) as
@@
which is an interior point of $(0,\infty)$. Then, according to Laplace’s method [53], the integral in the definition of EC can be expressed as
[…]
By substituting Eq. (164) into Eq. (16), we obtain Eq. (71).
@@ Appendix I Proof of Theorem 8
Similar to the proof of Lemma 2, the derivative of $g(x)$ is given by
$g^{\prime}(x)=\frac{-(\ln 2)^{-1}\left(\Xi+\Xi^{\prime}x\right)\left((1+\Xi x)\sqrt{(1+\Xi x)^{2}-1}-\sqrt{\frac{1}{N}}Q^{-1}(\epsilon)\right)}{(1+\Xi x)^{3}\sqrt{\big{(}1-(1+\Xi x)^{-2}\big{)}}},$ (165)
where for simplicity we omit the argument of $\Xi$. By comparing Eqs. (165) and (161), we find that if $\Xi+\Xi^{\prime}x\geq 0$, the remaining analysis is identical with the proof of Lemma 2.
@@
According to [28], the EC with the simple ARQ mechanism that has an unlimited number of retransmissions is given by
@@
Similar to the proofs of Lemma 2 and Theorem 8, if the power allocation scheme satisfies $\Xi(x)+\frac{{\rm d}\Xi(x)}{{\rm d}x}x\geq 0$, the integral $\int_{0}^{\infty}e^{-\theta NTr(x)}f(x){\rm d}x$ can be approximated as shown in Eq. (164) with the help of Laplace’s method. By substituting Eq. (164) into Eq. (166), we obtain Eq. (78).

### v4

@@ Appendix F Proof of Lemma 1
$\lim_{\gamma\to+\infty}\frac{1}{\sqrt{N}}\sqrt{\frac{-\log_{2}\epsilon(\gamma)}{\log_{2}^{2}\gamma}}$ (154a)
$=$ $\lim_{\gamma\to+\infty}\frac{\sqrt{-\log_{2}\epsilon(\gamma)}}{\sqrt{N}\log_{2}\gamma}$ (154b)
$=$ $\lim_{\gamma\to+\infty}\frac{\frac{\partial\sqrt{-\log_{2}\epsilon(\gamma)}}{\partial\log_{2}\gamma}}{\sqrt{N}+\frac{\partial\sqrt{N}}{\partial\log_{2}\gamma}\log_{2}\gamma}$ (154c)
@@
Eq. (154c) holds according to L’Hospital’s rule. As defined in Lemma 1, we let
$c_{4}=\lim_{\gamma\to+\infty}\frac{1}{1+\frac{\partial\sqrt{N}}{\partial\log_{2}\gamma}\frac{\log_{2}\gamma}{\sqrt{N}}}.$
Since $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$, we have
$\frac{\partial\sqrt{N}}{\partial\log_{2}\gamma}\frac{\log_{2}\gamma}{\sqrt{N}}\geq 0.$
Therefore, we conclude that $c_{3}\in[0,1]$.
Next, we let $Q^{-1}(\epsilon(\gamma))=x$, which can also be expressed as $\epsilon(\gamma)=Q(x)$. According to Eq. (103), we have
@@
By substituting $\epsilon(\gamma)$ for $v$ in Eq. (105), we obtain
$-\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}\leq x\leq\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}.$ (156)
Since $x>0$, we have $0<x\leq\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}$. Similarly, based on Eq. (103), we obtain Eq. (157), which is presented at the bottom of this page.
Based on the definition of $W(\cdot)$ and Eq. (157), we have
$x+\frac{1}{x}\geq\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}\text{ or }$ (158)
$x+\frac{1}{x}\leq-\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}.$
Since $x>0$, we will utilize $x+\frac{1}{x}\geq\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}$ in the following proof.
@@
Similarly, based on Eq. (158), we derive Eq. (160), which is shown at the bottom of this page. Eq. (160) holds because as $\gamma\to+\infty$, $Q^{-1}(\epsilon(\gamma))\to+\infty$. Therefore, by applying the squeeze theorem, we have completed the proof.
@@ Appendix G Proof of Corollary 3
Based on Lemma 1 in [50], we find that $\alpha_{A}(\theta)$ is an increasing function of $\theta$. Therefore, the inverse function of $\alpha_{A}(\cdot)$ exists, which is denoted by $\alpha^{-1}_{A}(\cdot)$. According to Eq. (20), the $\theta$ for the AWGN channel can be expressed as
@@
From Eq. (19), we know that $\chi\leq e^{-\theta L}$. Substituting Eq. (161) into $\chi\leq e^{-\theta L}\leq\chi_{\rm th}$, we have
@@
Since $\alpha_{A}(\theta)$ is an increasing function of $\theta$, it follows $R_{N}^{\epsilon}\geq\alpha_{A}\left(-\frac{1}{L}\ln\chi\right)$. Substituting Eq. (7) into this inequality, we obtain
$NC-\sqrt{NV}Q^{-1}(\epsilon)+\frac{\log_{2}N}{2}+$ $G(N,\gamma,\epsilon)$ (163)
$\geq\alpha_{A}\left(-\frac{1}{L}\ln\chi\right).$
From Eq. (163), we obtain Eq. (71).
@@ Appendix H Proof of Lemma 2
We put our focus on the expectation term in the definition of EC. For simplicity, let $g(x)=-r(x)$. The derivative of $g(x)$ is provided in Eq. (164), which is presented at the bottom of this page.
$g^{\prime}(x)=\frac{-(\ln 2)^{-1}\left(\Xi+\Xi^{\prime}x\right)\left((1+\Xi x)\sqrt{(1+\Xi x)^{2}-1}-\sqrt{\frac{1}{N}}Q^{-1}(\epsilon)\right)}{(1+\Xi x)^{3}\sqrt{\big{(}1-(1+\Xi x)^{-2}\big{)}}}.$ (168)
we find that $(1+\gamma x)\sqrt{(1+\gamma x)^{2}-1}$ is an increasing function of $x$, which equals 0 when $x=0$. Therefore, based on Eq. (164), there exists a stationary point $x^{*}$. When $x=x^{*}$, $g^{\prime}(x)=0$, $x<x^{*}$, $g^{\prime}(x)>0$, and $x>x^{*}$, $g^{\prime}(x)<0$. $x^{*}$ satisfies the following equation:
@@
The stationary point $x^{*}$ can be solved from Eq. (165). Specifically, we have
@@
which is an interior point of $(0,\infty)$. According to Laplace’s method [51], the integral in the definition of EC can be expressed as shown in Eq. (167), which is presented at the bottom of the last page. By substituting Eq. (167) into Eq. (16), we derive Eq. (72).
@@ Appendix I Proof of Theorem 8
Similar to the proof of Lemma 2, the derivative of $g(x)$ is given in Eq. (168), where we omit the argument of $\Xi$ for simplicity. Eq. (168) is presented at the bottom of this page. By comparing Eqs. (168) and (164), we find that if $\Xi+\Xi^{\prime}x\geq 0$, the remaining analysis is identical with the proof of Lemma 2. Therefore, we omit the details.
@@
According to [28], the EC with the simple ARQ mechanism that allows for unlimited retransmissions is given by
@@
Similar to the proofs of Lemma 2 and Theorem 8, if the power allocation scheme satisfies $\Xi(x)+\frac{{\rm d}\Xi(x)}{{\rm d}x}x\geq 0$, the integral $\int_{0}^{\infty}e^{-\theta NTr(x)}f(x){\rm d}x$ can be approximated as shown in Eq. (167) using Laplace’s method. By substituting Eq. (167) into Eq. (169), we obtain Eq. (79).
Acknowledgment
The authors would like to express their gratitude to the editor and anonymous reviewers for their careful reading and constructive comments. In particular, we appreciate their identification of the double-limit issue and their insightful suggestions on addressing it.